write the ordinary differential equation (1+sin y)dx = (2ycosy-x(secy+tan y))dy
1
Expert's answer
2020-03-26T14:11:43-0400

"(1+\sin y)x'+x\frac{1+\sin y}{\cos y}=2y\cos y\\
1) (1+\sin y)x'+x\frac{1+\sin y}{\cos y}=0\\
\frac{dx}{dy}=-\frac{x}{\cos y}\\
\frac{dx}{x}=-\frac{dy}{\cos y}\\
\ln|x|=-\ln|\tan(\frac{y}{2}+\frac{\pi}{4})|+\ln|c|\\
x_1=\frac{c}{\tan(\frac{y}{2}+\frac{\pi}{4})}\\
2) x=\frac{c(y)}{\tan(\frac{y}{2}+\frac{\pi}{4})}\\"

substitute x in equation

"(1+\sin y)\frac{c'(y)\tan(\frac{y}{2}+\frac{\pi}{4})-c(y)\frac{1}{2}\frac{1}{\cos^2(\frac{y}{2}+\frac{\pi}{4})}}{\tan^2(\frac{y}{2}+\frac{\pi}{4})}+\\
+\frac{c(y)}{\tan(\frac{y}{2}+\frac{\pi}{4})}\frac{1+\sin y}{\cos y}=2y\cos y\\
\frac{(1+\sin y)c'(y)}{\tan(\frac{y}{2}+\frac{\pi}{4})} -
\frac{(1+\sin y)c(y)}{2\tan^2(\frac{y}{2}+\frac{\pi}{4})\cos^2(\frac{y}{2}+\frac{\pi}{4})} +\\
+\frac{(1+\sin y)c(y)}{\tan(\frac{y}{2}+\frac{\pi}{4})\cos y}=2y\cos y\\
\frac{(1+\sin y)c'(y)}{\tan(\frac{y}{2}+\frac{\pi}{4})} -\frac{(1+\sin y)c(y)}{2\sin^2(\frac{y}{2}+\frac{\pi}{4})}+\\
+\frac{(1+\sin y)c(y)}{\tan(\frac{y}{2}+\frac{\pi}{4})\cos y}=2y\cos y"

show that

"2\sin^2(\frac{y}{2}+\frac{\pi}{4})=2\tan(\frac{y}{2}+\frac{\pi}{4}) \cos y\\
2\sin(\frac{y}{2}+\frac{\pi}{4})=\frac{\cos y}{\cos(\frac{y}{2}+\frac{\pi}{4})}\\
2\sin(\frac{y}{2}+\frac{\pi}{4})\cos(\frac{y}{2}+\frac{\pi}{4})=\cos y\\
\sin2(\frac{y}{2}+\frac{\pi}{4})=\cos y\\
\sin(y+\frac{\pi}{2})=\cos y\\
\cos y=\cos y"

so

"\frac{(1+\sin y)c'(y)}{\tan(\frac{y}{2}+\frac{\pi}{4})} =2y\cos y\\
c'(y)=\frac{2y\cos y\tan(\frac{y}{2}+\frac{\pi}{4})}{1+\sin y} =\\
=\frac{2y\cos y\tan(\frac{y}{2}+\frac{\pi}{4})}{\cos^2\frac{y}{2}+2\cos\frac{y}{2}\sin\frac{y}{2}+
\sin^2\frac{y}{2}} =\\
=\frac{2y(\cos^2\frac{y}{2}-\sin^2\frac{y}{2})\tan(\frac{y}{2}+\frac{\pi}{4})}{(\cos\frac{y}{2}+\sin\frac{y}{2})^2} =\\
=\frac{2y(\cos\frac{y}{2}-\sin\frac{y}{2})\sin(\frac{y}{2}+\frac{\pi}{4})}{\cos(\frac{y}{2}+\frac{\pi}{4})(\cos\frac{y}{2}+\sin\frac{y}{2})} =\\
=\frac{2y(\cos\frac{y}{2}-\sin\frac{y}{2})(\sin\frac{y}{2}\cdot\frac{\sqrt{2}}{2}+\cos\frac{y}{2}\cdot\frac{\sqrt{2}}{2})}{(\cos\frac{y}{2}\cdot\frac{\sqrt{2}}{2}-\sin\frac{y}{2}\cdot\frac{\sqrt{2}}{2})(\cos\frac{y}{2}+\sin\frac{y}{2})} =2y\\
c'(y)=2y\\
c(y)=y^2+c_1"

So solution of equation is

"x=\frac{y^2+c_1}{\tan(\frac{y}{2}+\frac{\pi}{4})}"